Explain how the banking money multiplier works? (Use words as well as math for a complete answer)
Answer #1

The formula for the money multiplier is as follows:-

1/R

Where R is the reserve requirement

Suppose the reserve requirement is 10

Then the money multiplier = 1/0.10 = 10

And if the fed buys $1000, then the change in the money supply would be 1000*10 = 10000